OpenStudy (anonymous):

\[X:\sqrt{3x} +9 =-6\]

OpenStudy (turingtest):

1) subtract 9 from both sides 2) square both sides 3) divide both sides by 3

OpenStudy (anonymous):

okay hold on I will try it

OpenStudy (turingtest):

let me know if/where you get stuck

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I dont think there is a solution

OpenStudy (turingtest):

hm... could you show your steps? what do you have after the first step?

OpenStudy (anonymous):

thanks hold on im really bad at math great at english lol

OpenStudy (anonymous):

i got 9

OpenStudy (turingtest):

for the final answer? no, that's not right I don't think just subtract 9 from both sides first what's left?

OpenStudy (anonymous):

3

OpenStudy (turingtest):

where did x go?\[\sqrt{3x}+9=-6\]subtracting 9 from both sides gives\[\sqrt{3x}+\cancel9^0-\cancel9^0=-6-9\]\[\sqrt{3x}=-15\]if this is novel to you I think you need to step back to the basics of doing arithmetic on each side of the euation the next step is to square both sides remember that\[(\sqrt x)^2=x\]and don't forget tot square \(both\) sides

OpenStudy (anonymous):

the choices I have for answers are 9, 1/9, 81, no real solution